OpenStudy (anonymous):

What is the volume of a sphere with a radius of 6 inches? 112Ï€ cubic inches 288Ï€ cubic inches 290Ï€ cubic inches 300Ï€ cubic inches

OpenStudy (anonymous):

Use the formula:\[Volume = \dfrac{4}{3} \pi r^{3}\]Where \(r\) is the radius.

OpenStudy (anonymous):

@Algorithmic someone suggested that earlier i still cant find the answer :(

OpenStudy (anonymous):

Okay, let us go through the steps together. Substitute \(6\) in place of \(r\). Do you know how to do that?

OpenStudy (anonymous):

@Algorithmic my assignment is going to time out in lie 30 seconds can we speed this up

OpenStudy (anonymous):

\[Volume = \dfrac{4}{3} \pi 6^{3}\] So now that we have substituted \(r\) with \(6\). (If you do not follow just ask.) Now just multiply everything in regards to BODMAS/PEMDAS/BEDMAS, and you should get your volume.

OpenStudy (anonymous):

Lol @Algorithmic so whats the answer?

OpenStudy (zale101):

Use the information what @Algorithmic gave you and try to do the problem. Don't forget to show your work!

OpenStudy (anonymous):

Omg you guys why cant u juss say the answer im in a rush

OpenStudy (zale101):

Unfortunately. The goal of openstudy is to assist you with solving the problem, not by giving the answer. We assist YOU to solve it, we don't assist you by copying our answers.

OpenStudy (zale101):

You are not learning when you copy answers, and that is not the goal of openstudy and your school.